PESSTO Year 3+4 - Final SOFI spectra reduction


Summary:
SOFI spectra reduced with PESSTO pipeline version 2.2.9

109 SOFI spectra reduced (64 BG + 45 RG).
6 SOFI spectra not reduced (3 BG + 3 RG).
8 SOFI spectra of the 109 reduced spectra (~7%) have a very low S/N (4 BG + 4 RG). Subjective estimate.


2014-09-03:
ASASSN-14gm BG+RG spectra have no accompanying telluric standard observed. Used a telluric standard HIP3086 (airmass 1.25) from the night 2014-09-22 to reduce ASASSN-14gm (airmass ranging from 1.4 to 1.1).

The second ASASSN-14gm sequence fine. No proper trace in the first sequence:
SOFI.2014-09-04T02:53:02.309.fits
SOFI.2014-09-04T02:58:47.764.fits
SOFI.2014-09-04T03:04:29.940.fits
SOFI.2014-09-04T03:10:21.952.fits
SOFI.2014-09-04T03:17:24.604.fits
SOFI.2014-09-04T03:26:20.939.fits
SOFI.2014-09-04T03:35:07.439.fits
SOFI.2014-09-04T03:44:00.494.fits
Used only the second sequnce.

Feige110 flux standard not recognised as such by the pipeline. Used a spectroscopic standard LTT3218 from the night 2014-09-29.

2014-09-22:
No flux standard observed. Used a spectroscopic standard LTT3218 from the night 2014-09-29.

ASASSN-14ha has no visible trace in the last two frames
SOFI.2014-09-23T05:21:53.657.fits
SOFI.2014-09-23T05:30:46.712.fits
of a four exposure RG sequence. Used only the first two exposures.

ASASSN-14ha BG+RG spectra have no accompanying telluric standard observed. Telescope closed due to clouds. Used HIP3086 which is a telluric standard for ASASSN-14gm from the same night.

2014-10-13:
ASASSN-14il BG spectrum has no accompanying telluric standard observed. Used HIP020378 which is a telluric standard for ASASSN-14ha from the same night.

2014-10-23:
HIP3990 was likely aimed to be observed as the telluric standard for ASASSN-14il after the science observations. HIP3990 has been reported in the nightlog, however in reality HIP110512 was observed (with very poor airmass). Pipeline used for the telluric correction of ASASSN-14il an earlier observation of HIP110512 which was observed as an alternative calibration for ASASSN-14jb and has a more similar airmass to ASASSN-14il.

2014-11-14:
Feige110 observations reported to have slit problems in the nightlog. Using only the last BG+RG sequence.

ASASSN-14jb BG 12 observation sequence aborted after frame 9/12 due to a lost guide star. Problems with the 2d reductions. Used only the first 6 exposures.

2014-11-23:
ASASSN-14jb BG sequence has no dither between 'B' and 'A' exposures
SOFI.2014-11-24T00:52:00.656.fits
SOFI.2014-11-24T00:59:05.429.fits
of an ABBA sequence. These two frames were not used.
As a byproduct of normally continuing dither sequence, the following ASASSN-14jb frames have the trace close to the right side of the FOV:
SOFI.2014-11-24T01:13:24.815.fits
SOFI.2014-11-24T01:20:26.310.fits
SOFI.2014-11-24T01:41:50.469.fits
SOFI.2014-11-24T01:48:51.967.fits
However, the target ASASSN-14jb is on the slit and the data can be used.

2014-12-21:
PESSTOESO154-G10 BG spectrum pipeline reduction is using HIP003559 as the telluric standard by default. This is a telluric standard observed for ASASSN-14il. Forced the pipeline to use HIP013315 which was observed after PESSTOESO154-G10 (named incorrectly as HIP013515 in the nightlog).

2015-01-19:
No flux standard observed. Used a spectroscopic standard GD71 from the night 2014-12-21.

LSQ14gqk BG spectrum pipeline reduction is using HIP10694 as the telluric standard by default. Forced the pipeline to use HIP27185 which was observed after LSQ14gqk.

ASASSN-14lw BG final spectrum has a very low S/N.

2015-02-15:
PSNJ10544216 RG final spectrum has a very low S/N.

2015-02-24:
HIP056242 BG telluric standard observed before PS15ae is causing a peculiar continuum break in the PS15ae BG spectrum. Forced the pipeline to use instead HIP056614 observed after PS15ae.

PS15ae RG spectrum OB aborted at 4/8 exposures. Final spectrum has a very low S/N.

PSNJ10544216-2104138 BG sequence reduction resulted in a noisy upper right quadrant in the 2d image pair subtraction from the original frames
SOFI.2015-02-25T05:59:35.877.fits
SOFI.2015-02-25T06:12:17.217.fits

2015-03-10:
PSNJ10544216-2104138 BG spectrum pipeline reduction is using HIP077338 as the telluric standard by default. This is a telluric standard observed for PS15ae. Forced the pipeline to use HIP090869 which was observed after PSNJ10544216-2104138.

2015-03-11:
No spectroscopic standard observed. Using EG274 from the previous night 2015-03-10.

2015-03-12:
PSNJ07361576-6930230 BG spectrum dither sequence has no offset between the two first frames
SOFI.2015-03-13T00:05:24.908.fits
SOFI.2015-03-13T00:11:07.084.fits
These frames were not used.

PSNJ07361576-6930230 RG sequence reduction resulted in a noisy upper right quadrant in the 2d image pair subtraction from the original frames
SOFI.2015-03-13T00:29:50.796.fits
SOFI.2015-03-13T00:38:47.131.fits
However, this is not affecting the trace of the SN.

2015-03-13:
No spectroscopic standard observed. Using LTT3218 from the previous night 2015-03-12.

PSNJ07361576-6930230 BG+RG frames have a noisy upper right quadrant in the all the 2d image pair subtractions. This is affecting the trace of the SN in 2/4 BG and 2/4 RG frames.

2015-03-18:
No spectroscopic standard observed. Using LTT7987 from the night 2015-03-20.

PSNJ07361576-6930230 BG+RG spectroscopic sequence does not dither between the exposures (Furthermore, noisy upper right quadrant in all the 2d image pair subtractions). DATA NOT REDUCED!!!

2015-03-20:
Serious noise problems with the SOFI detector reported in the nightlog.

HIP029601 telluric standard BG+RG frames observed for PS15ae. However, data useless due to the detector noise. These frames were not used. Used HIP054366 which is a telluric standard for PSNJ07361576-6930230.

PSNJ07361576-6930230 BG spectrum sequence is observed with very large dither steps? From the four exposures these two are not useful as the trace is outside the FOV
SOFI.2015-03-21T01:18:46.510.fits
SOFI.2015-03-21T01:30:14.138.fits
These frames were not used.

PSNJ07361576-6930230 RG final spectrum has a very low S/N.

PS15br BG frames do not seem to have any signal. DATA NOT REDUCED!!!

2015-04-09:
SNhunt281 BG+RG spectroscopic sequence does not dither between the exposures. DATA NOT REDUCED!!!

SN2015F RG spectrum dither sequence has no offset between the two first frames
SOFI.2015-04-10T02:19:42.504.fits
SOFI.2015-04-10T02:31:14.008.fits
These frames were not used.

LTT3218 flux standard BG+RG spectroscopic sequence does not dither between exposures. Using spectroscopic standard LTT3128 from the night 2015-04-11.

2015-04-11:
SN2015F BG spectrum dither sequence has no offset between the two last frames
SOFI.2015-04-12T01:52:21.556.fits
SOFI.2015-04-12T02:00:03.737.fits
These frames were not used.
As a byproduct of normally continuing dither sequence, the following SN2015F frames have the trace close to the right side of the FOV:
SOFI.2015-04-12T02:20:40.579.fits
SOFI.2015-04-12T02:32:12.081.fits
SOFI.2015-04-12T02:43:50.142.fits
However, the target SN2015F is on the slit and the data can be used.

ASASSN-15ga BG spectrum pipeline reduction is using HIP040794 as the telluric standard by default. This is a telluric standard observed for SN2015F. Forced the pipeline to use HIP66974 which was observed after SN2015F.

2015-04-18:
SNhunt281 RG spectra have no signal. Instrumental problem reported in the nightlog. DATA NOT REDUCED!!! 

LSQ15adm BG observing sequence was likely aborted after 2 exposures. Data reduced nonetheless. Final spectrum has a very low S/N ratio.

2015-04-27:
SN2015F BG sequence of 8 exposures has no dither between the 7 first exposures. Using only the last two frames.

2015-08-14:
No spectroscopic standard observed. Using EG274 from the night 2015-10-04.

PSNJ23002463+0137368 RG spectral sequence not finished, only 5/8 exposures. Final spectrum has a very low S/N.

2015-09-04:
Feige110 spectroscopic standard BG+RG spectrum dither sequence has no offset between the exposures. Using EG274 from the night 2015-10-04.

HIP094926 telluric standard BG spectrum has a peculiar continuum shape. Likely something wrong in the exposure. Resulting ASASSN-15oz BG spectrum reduced using HIP094926 has also a peculiar continuum.

2015-10-04:
PSNJ02051332+0606084 BG spectrum pipeline reduction is using HIP102021 as the telluric standard by default. This is a telluric standard observed for ASASSN-15oi. Forced the pipeline to use HIP024336 which was observed after PSNJ02051332+0606084.

ASASSN-15oi BG final spectrum has a low S/N.

2016-03-08:
No spectroscopic standard observed. Using EG274 from the previous night 2016-03-07.

Gaia16afe BG final spectrum has a low S/N.
Comments